OpenStudy (anonymous):

Can someone help me work through this problem please , Having trouble understanding . Its Systems of equations.

OpenStudy (anonymous):

what is the problem?

OpenStudy (anonymous):

4. Follow the 5 steps below to complete this problem. (4 points) Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: • Drive a boat • Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

its a 5 step equation

OpenStudy (anonymous):

good lord, okay i guess we can try this

OpenStudy (anonymous):

haha, you can see my struggle !

OpenStudy (anonymous):

how about i learn how to read. you are not supposed to pick the speed of the boat, you are supposed to pick the current how about 4 mph for the current?

OpenStudy (anonymous):

Sounds good .

OpenStudy (anonymous):

so 4 mph for the current

OpenStudy (anonymous):

so lets start with 1) justin drives a boat upstream and back 2) the current is 4 mph 3) it takes him 5 hours to go up steam 4) it takes him 4 hours to come back

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

Very!

OpenStudy (anonymous):

so now we have to solve it

OpenStudy (anonymous):

Thats where i have the trouble at

OpenStudy (anonymous):

damn okay we can do that too. i suppose the question is a) how far did he go or? b) what is the speed of the boat in still water?

OpenStudy (anonymous):

so lets call the rate in still water \(r\) for "rate" and the distance he travelled \(D\) for distance. since the current is 4 mph the rate with the current is \(r+4\) and the rate against the current is \(r-4\) so far so good?

OpenStudy (anonymous):

now we can use "distance equals rate times time" the rate with the current is \(r+4\) and the time with the current is 4 hours the rate against the current is \(r-4\) and the time against the current is 5 hours since the distance going is evidently the same as the distance returning, you know that \[D=(r-4)\times 5=(r+4)\times 4\]

OpenStudy (anonymous):

in other words, your last job is to solve \[5(r-4)=4(r+4)\] for \(r\) that is the last step

OpenStudy (anonymous):

so you just distribute 5(r-4)= 4(r+4) 5r-20 =4r+16?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

add the common factors 5r-20 =4r+16? 5r 5r 20= r+16 ? 16 16 4 =r

OpenStudy (anonymous):

last steps not so good lets start with \[5r-20=4r+16\] and subtract \(4r\) from both sides

OpenStudy (anonymous):

ok so 5r−20=4r+16 4r 4r r -20=16?

OpenStudy (anonymous):

yes, better now add 20 to both sides

OpenStudy (anonymous):

r-20=16 20 20 r= 36

OpenStudy (anonymous):

yes, good on your first attempt you actually made two mistakes \(4r-5r=-r\) not \(r\) and also you dropped the minus sign in front of the 20

OpenStudy (anonymous):

so if you had wanted to solve by subtracting \(5r\) from both sides it would have looked like \[4r-20=4r+16\] \[-20=-r+16\] \[-36=-r\] \[36=r\]

OpenStudy (anonymous):

I get my negatives and postives mixed up im pretty bad with that. /: thanks for the help ! i appreciate it !

OpenStudy (anonymous):

yw, good luck!!

OpenStudy (anonymous):

could you help me with one more?